Find the future value of an investment of $3,900 made today for the following rates and periods: (If you solve this problem with algebra round intermediate calculations to 4 decimal places, in all cases round your answers to the nearest penny.)

a. 6.25 percent compounded semiannually for 12 years. Future value?

b. 7.63 percent compounded quarterly for 6 years. Future value?

c. 8.9 percent compounded monthly for 10 years. Future value?

d. 10 percent compounded daily for 3 years. Future value?

e. 8 percent compounded continuously for 2 years. Future value?
Answer #1

Future Value=Present Value*(1+r/m)^(m*t)

1.
=3900*(1+6.25%/2)^(2*12)=8162.05814755269

2.
=3900*(1+7.63%/4)^(4*6)=6137.79174974346

3.
=3900*(1+8.9%/12)^(12*10)=9465.8668472022

4.
=3900*(1+10%/365)^(3*365)=5264.2330462635

5.
=3900*e^(0.08*2)=$4,576.692